# Rank-One Modified Value Iteration

### v1

@@ 5. Numerical simulations
In this section, we compare the performance of several planning and learning algorithms with our proposed methods. The experiments are conducted on Garnet [2] and Graph MDPs [7], focusing on the Bellman errors $\left\|\mathbf{T}(\bm{v}_{k})-\bm{v}_{k}\right\|_{\infty}$ and $\left\|\overline{\mathbf{T}}(\bm{q}_{k})-\bm{q}_{k}\right\|_{\infty}$ and the value errors $\left\|\bm{v}_{k}-\bm{v}^{\star}\right\|_{\infty}$ and $\left\|\bm{q}_{k}-\bm{q}^{\star}\right\|_{\infty}$. In all of our experiments, we run policy iteration (PI) until it converges to calculate the optimal values $\bm{v}^{\star}$ and $\bm{q}^{\star}$ for reference. Garnet and Graph MDPs are particularly compelling for our empirical analysis as we can run PI to compute the optimal values, enabling us to measure value errors throughout the iterations. The Garnet MDPs have a state size of $n=200$, an action size of $m=5$, and randomly generated transition probabilities and costs with a branching factor of 10. For our numerical experiments, we consider 25 randomly generated instances of Garnet MDPs and report the three quantiles of the errors. For the Graph MDPs, we use the same configuration as described in [7], providing a complementary benchmark to validate the effectiveness of our proposed method. We also note that the supplementary material includes the Python package implementing the algorithms and the experiments discussed in this section. In what follows, we report the result of our simulations for the planning and learning problems.
@@ 6. Limitations and future research directions
Let us start by noting that the provided theoretical results in Theorems 3.3 and 4.2 guarantee the convergence of the proposed algorithms with the same rate as the standard VI and QL algorithms. However, our numerical experiments with Garnet and Graph MDPs in Section 5 show that the proposed algorithms have a faster convergence rate compared to standard VI and QL and their accelerated versions. This gap can be explained by the fact that our proof technique does not exploit that the vectors $\bm{d}_{k}$ and $\widehat{\bm{d}}_{k}$ used in the update rule are specifically constructed to approximate the stationary distribution of the Markov chain induced by the greedy policy (see Appendices A.3 and A.4 for details). That is, the convergence of R1-VI and R1-QL algorithms is guaranteed for any choice of $\bm{d}_{k}\in\Delta(\mathcal{S})$ and $\widehat{\bm{d}}_{k}\in\Delta(\mathcal{S}\times\mathcal{A})$. In fact, when the stationary distribution concentrates on a single state, as happens when there is an absorbing state with zero reward, the second term in the R1VI update (Equation (12)) vanishes. Appendix B.3 provides an empirical analysis in Gridworld [30], which includes an absorbing state and thus violates the assumption of Lemma 3.2. Moreover, the provided proof of convergence shows that the greedy policies generated with respect to the iterates of R1-VI and R1-QL are the same as those for the standard VI and QL algorithms, respectively (see Lemmas A.2 and A.4). In other words, the proposed algorithms do not affect the speed of convergence to the optimal policy compared to VI and QL. Nevertheless, at least in the case of R1-VI, the faster convergence in the value spaces leads to a faster termination of the algorithm for a given performance bound for the greedy policy. In this regard, let us also note that the mismatch between convergence in value space and policy space also arises in other “accelerated” VI/QL algorithms; see Appendix B.4.
@@
Third, we note that the proposed R1-QL algorithm 2 is a synchronous algorithm that updates all state-action pairs $(s,a)\in\mathcal{S}\times\mathcal{A}$ of the Q-function $\bm{q}_{k}$ at each iteration $k$. This algorithm can be modified in a standard fashion for the asynchronous case. However, the provided convergence analysis can not be extended for the corresponding asynchronous algorithm in a straightforward manner. Moreover, the straightforward asynchronous implementation of R1-QL leads to an $\operatorname{\mathcal{O}}(nm)$ per-iteration complexity for updating a single entry $(s,a)\in\mathcal{S}\times\mathcal{A}$ at each iteration $k$, which is higher than the $\operatorname{\mathcal{O}}(m)$ per-iteration complexity of the standard asynchronous QL algorithm. We note that the Zap Q-leaning algorithm [7] also suffers from this issue. Addressing these issues requires a more involved analysis and modification of the proposed algorithm in the asynchronous case, which we leave for future research.
@@ B.3. Reducible MDPs
In Absorbing-Gridworld (left side of Figure 7), R1-VI yields the lowest value and Bellman error, apart from PI, across all $\gamma$ values. However, in Terminal Gridworld (right side of Figure 7), R1-VI performs slightly worse than the other accelerated algorithms. This is explained by the fact that under the optimal policy, the stationary distribution concentrates on the absorbing state, which provides zero reward and hence zero Bellman error when the values are initialized to zero. Consequently, the second term in the R1-VI update (Equation (9)) vanishes. In contrast, in Absorbing Gridworld, the Bellman error at the absorbing state is not immediately zero, hence the second term in R1-VI contributes to improve convergence.
@@ B.4. Policy performance
In both Figures 9 and 8, the proposed R1VI and R1QL algorithms match the policy performance of VI and QL, respectively, across all iterations. Moreover, VI in the planning setting and QL in the learning setting produce among the highest policy performance observed across both MDPs, despite exhibiting the slowest convergence in the value space.
Acknowledgements
The authors would like to thank the reviewers for their useful comments.

### v2

@@ 5. Numerical simulations
In this section, we compare the performance of several planning and learning algorithms with our proposed methods. The experiments are conducted on Garnet [2] and Graph MDPs [7], focusing on the Bellman errors $\left\|\mathbf{T}(\bm{v}_{k})-\bm{v}_{k}\right\|_{\infty}$ and $\left\|\overline{\mathbf{T}}(\bm{q}_{k})-\bm{q}_{k}\right\|_{\infty}$ and the value errors $\left\|\bm{v}_{k}-\bm{v}^{\star}\right\|_{\infty}$ and $\left\|\bm{q}_{k}-\bm{q}^{\star}\right\|_{\infty}$. In all of our experiments, we run policy iteration (PI) until it converges to calculate the optimal values $\bm{v}^{\star}$ and $\bm{q}^{\star}$ for reference. Garnet and Graph MDPs are particularly compelling for our empirical analysis as we can run PI to compute the optimal values, enabling us to measure value errors throughout the iterations. The Garnet MDPs have a state size of $n=200$, an action size of $m=5$, and randomly generated transition probabilities and costs with a branching factor of 10. For our numerical experiments, we consider 25 randomly generated instances of Garnet MDPs and report the three quantiles of the errors. For the Graph MDPs, we use the same configuration as described in [7], providing a complementary benchmark to validate the effectiveness of our proposed method. In what follows, we report the result of our simulations for the planning and learning problems.
@@ 6. Limitations and future research directions
Let us start by noting that the provided theoretical results in Theorems 3.3 and 4.2 guarantee the convergence of the proposed algorithms with the same rate as the standard VI and QL algorithms. However, our numerical experiments with Garnet and Graph MDPs in Section 5 show that the proposed algorithms have a faster convergence rate compared to standard VI and QL and their accelerated versions. This gap can be explained by the fact that our proof technique does not exploit that the vectors $\bm{d}_{k}$ and $\widehat{\bm{d}}_{k}$ used in the update rule are specifically constructed to approximate the stationary distribution of the Markov chain induced by the greedy policy (see Appendices A.3 and A.4 for details). That is, the convergence of R1-VI and R1-QL algorithms is guaranteed for any choice of $\bm{d}_{k}\in\Delta(\mathcal{S})$ and $\widehat{\bm{d}}_{k}\in\Delta(\mathcal{S}\times\mathcal{A})$. In fact, when the stationary distribution concentrates on a single state, as happens when there is an absorbing state with zero reward, the second term in the R1VI update rule (12) vanishes. Appendix B.3 provides an empirical analysis in Gridworld [30], which includes an absorbing state and thus violates the assumption of Lemma 3.2. Moreover, the provided proof of convergence shows that the greedy policies generated with respect to the iterates of R1-VI and R1-QL are the same as those for the standard VI and QL algorithms, respectively (see Lemmas A.2 and A.4). In other words, the proposed algorithms do not affect the speed of convergence to the optimal policy compared to VI and QL. Nevertheless, at least in the case of R1-VI, the faster convergence in the value spaces leads to a faster termination of the algorithm for a given performance bound for the greedy policy. In this regard, let us also note that the mismatch between convergence in value space and policy space also arises in other “accelerated” VI/QL algorithms; see Appendix B.4.
@@
Third, we note that the proposed R1-QL algorithm 2 is a synchronous algorithm that updates all state-action pairs $(s,a)\in\mathcal{S}\times\mathcal{A}$ of the Q-function $\bm{q}_{k}$ at each iteration $k$. This algorithm can be modified in a standard fashion for the asynchronous case. However, the provided convergence analysis can not be extended for the corresponding asynchronous algorithm in a straightforward manner. Moreover, the straightforward asynchronous implementation of R1-QL leads to an $\operatorname{\mathcal{O}}(nm)$ per-iteration complexity for updating a single component $(s,a)\in\mathcal{S}\times\mathcal{A}$ at each iteration $k$, which is higher than the $\operatorname{\mathcal{O}}(m)$ per-iteration complexity of the standard asynchronous QL algorithm. We note that the Zap Q-leaning algorithm [7] also suffers from this issue. Addressing these issues requires a more involved analysis and modification of the proposed algorithm in the asynchronous case, which we leave for future research.
@@ B.3. Reducible MDPs
In Absorbing-Gridworld (left side of Figure 7), R1-VI yields the lowest value and Bellman error, apart from PI, across all $\gamma$ values. However, in Terminal Gridworld (right side of Figure 7), R1-VI performs slightly worse than the other accelerated algorithms. This is explained by the fact that under the optimal policy, the stationary distribution concentrates on the absorbing state, which provides zero reward and hence zero Bellman error when the values are initialized to zero. Consequently, the second term in the R1-VI update rule (9) vanishes. In contrast, in Absorbing Gridworld, the Bellman error at the absorbing state is not immediately zero, hence the second term in R1-VI contributes to improve convergence.
@@ B.4. Policy performance
In both Figures 8 and 9, the proposed R1VI and R1QL algorithms match the policy performance of VI and QL, respectively, across all iterations. Moreover, VI in the planning setting and QL in the learning setting produce among the highest policy performance observed across both MDPs, despite exhibiting the slowest convergence in the value space.